OpenStudy (anonymous):

Given f(x)=5x+5 and g(x)=6x+1 then what is f(g(−8)) ? Given f(x)=10x+8 and g(x)=1x+4 then what is (f+g)(3)) ? Given f(x)=3x+4 and g(x)=8x+1 then what is (f−g)(0)) ?

OpenStudy (anonymous):

Given f(x)=5x+5 and g(x)=6x+1, then what is f(g(−8)) ? First, substitute -8 to g(x). \(g(x)=6x+1\rightarrow g(x)=6(-8)+1\). Solve for x.

OpenStudy (anonymous):

Tell me what x is, so can move to step two.

OpenStudy (anonymous):

x = -47

OpenStudy (anonymous):

Good! Now, we need to substitute g(x)=-47 to f(x)=5x+5. \[f(x)=5x+5\rightarrow f(g(-8))= 5(-47)+5\]

OpenStudy (anonymous):

Solve it and that's the answer. :)

OpenStudy (anonymous):

-230?

OpenStudy (anonymous):

Correct. :)

OpenStudy (anonymous):

Can you help me with the rest ?

OpenStudy (anonymous):

Given f(x)=10x+8 and g(x)=1x+4 then what is (f+g)(3)) ? this is 45? Given f(x)=3x+4 and g(x)=8x+1 then what is (f−g)(0)) ? this is -1?

OpenStudy (anonymous):

Given f(x)=10x+8 and g(x)=1x+4 then what is (f+g)(3)) ? Add f(x) and g(x) first.

OpenStudy (anonymous):

11x + 12

OpenStudy (anonymous):

Then substitute 3 to x.

OpenStudy (anonymous):

33 + 12 = 45

OpenStudy (anonymous):

Oh wait, is that your answer? 45?

OpenStudy (anonymous):

lol yes

OpenStudy (anonymous):

I didn't see it at first. :D Yes you're right! I'll check your other answer. :)

OpenStudy (anonymous):

-5x - 3 = -5(0) -3 = -3

OpenStudy (anonymous):

\[3x+4-8x-1\]

OpenStudy (anonymous):

\[-5x+3\] Substitute 0 to x.

OpenStudy (anonymous):

-5x - 3 = -5(0) -3 = -3

OpenStudy (anonymous):

Wrong. \(-5x+3\) Substitute 0 to x.

OpenStudy (anonymous):

no positive 3

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Thank you can you help me with something else .

OpenStudy (anonymous):

Just post another question and I'd be happy to help. :)

OpenStudy (anonymous):

okay . thanks . !!!